# TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models

### v1

@@ 4.1 Models
We evaluate several popular LLMs, including both open-source and closed-source models, with parameter sizes ranging from 6B to 175B [^1] Since OPENAI does not disclose ChatGPT-3.5/4’s size publicly, we use the size of GPT-3, which is 175 billion.. Unless otherwise stated, we use aligned models in the following experiments. By default, we use 13B model for LLaMA2/Vicuna-1.5/Baichuan2 and 11B model for FLAN-T5.
ChatGPT-3.5/GPT-4 (Ouyang et al., 2022; OpenAI, 2023)
ChatGPT is a chat model aligned through SFT and RLHF based on GPT-3 (Brown et al., 2020). GPT-4 is an upgraded version of ChatGPT with enhanced reasoning capabilities, making it the most powerful LLM. Unless otherwise stated, ChatGPT refers to gpt-3.5-turbo-0613 and GPT-4 refers to gpt-4-0613.
Llama2/Vicuna-1.5 (Touvron et al., 2023; Chiang et al., 2023)
LLaMA2 is an open foundation model trained on 2T tokens with efficient grouped-query attention Ainslie et al. (2023). LLaMA2-chat is the official aligned model with SFT and RLHF, and Vicuna-1.5 is an unofficial aligned model with SFT only.
Baichuan2 (Yang et al., 2023)
is an open foundation model pre-trained on 2.6T tokens and outperforms LLaMA2 on several benchmarks. Baichuan2-chat is the official aligned model with SFT and RLHF.
Mistral (Jiang et al., 2023)
is a 7B open foundation model incorporating efficient grouped-query attention (Ainslie et al., 2023) and sliding windows attention (Beltagy et al., 2020). It achieves the strongest performance among models of its size, even surpassing LLaMA2-13B. Mistral-instruct is the officially aligned model with SFT only.
ChatGLM3 (Zeng et al., 2023)
is an open-source bilingual model for Chinese and English, exhibiting competitive performance in models with sizes under 10B.
FLAN-T5 (Shen et al., 2023)
is an open-source instruction-following model built on top of T5 (Raffel et al., 2020) through instruction fine-tuning.
@@ 4.2 Implementation Details
We access closed-source models through OpenAI API. For open-source models, we deploy them locally and then access them through FastAPI. We set the temperature to 0 for greedy search in all experiments. In zero-shot-cot, we use Let’s think step by step as the reasoning trigger. To improve answer extraction accuracy, we use answer trigger Therefore, the answer is to prompt models to summarize the answer. In few-shot scenarios, we manually create 2-4 instances, varying with datasets. For more details about demonstrations and instructions, please refer to Appendix A.4.
Method Symbolic Commonsense Event Temporal Overall
TimeXNLI Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE
s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM I-EM S-EM O-EM C-EM L2-EM L3-EM Acc Sym. Comm. Event Overall
Human 98 96 92 100.0 64 80.8 75.8 87.1 97.8 - 100 89 87 82 84 76 96 94 82.5 96.5 84.3 88.0 88.5
GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 40.4 44.4 49.0 22.0 91.0 94.0 64.8 75.8 65.4 59.3 63.5
  + CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 33.0 43.4 53.0 20.0 93.0 93.0 58.0 77.0 66.2 58.3 63.2
  + FS 85.3 73.3 53.3 100.0 51.0 64.8 77.5 88.3 85.0 94.6 88.6 59.2 40.0 59.6 48.0 25.3 86.0 94.8 62.8 78.0 78.6 63.0 69.5
  + FS/CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 44.4 48.5 44.0 22.0 91.0 93.0 66.4 85.0 68.3 60.8 66.5
GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.5 39.2 69.1 62.3 60.5 29.5 36.5 37.5 21.0 73.6 61.8 57.4 60.3 49.0 49.8 51.2
  + CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 29.0 35.8 38.5 24.0 32.0 54.2 52.0 50.8 35.1 43.5 42.6
  + FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 37.9 37.8 43.5 16.0 77.7 70.0 55.0 53.9 61.2 52.1 54.8
  + FS/CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 37.5 38.1 37.5 33.0 86.2 68.0 50.2 61.1 40.2 54.0 51.8
Baichuan2 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 29.0 31.0 34.0 30.0 40.0 45.0 49.0 37.9 37.5 42.9 40.7
  + CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 36.0 39.0 27.0 29.0 46.0 46.0 55.0 34.5 33.5 43.8 39.8
  + FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 35.0 37.0 31.0 19.0 73.0 48.0 48.0 46.1 46.4 47.3 46.9
  + FS/CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 27.0 38.0 34.0 33.0 72.8 43.0 44.0 48.5 36.8 46.9 44.7
LLaMA2 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 21.0 28.0 23.0 18.0 43.0 55.0 49.0 33.9 27.8 40.5 36.1
  + CoT 36.0 50.0 38.0 6.0 7.3 39.1 14.0 51.7 10.0 36.9 - 45.0 30.0 20.0 18.0 21.0 43.0 56.0 47.0 32.5 26.5 40.7 36.0
  + FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.5 15.0 62.3 40.2 24.0 17.0 11.0 5.0 22.0 54.0 50.0 47.0 45.1 35.4 32.6 35.3
  + FS/CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 21.0 34.0 23.0 7.0 72.0 54.0 50.0 43.8 28.4 40.9 38.4
Vicuna-1.5 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 29.0 38.0 22.0 17.0 14.0 13.0 46.0 34.0 31.5 34.2 33.4
  + CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 31.0 16.0 25.0 13.0 31.0 29.0 51.0 33.3 22.2 35.2 31.8
  + FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 23.0 38.0 26.0 18.0 51.0 28.0 56.0 43.4 31.1 38.5 37.3
  + FS/CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 27.0 39.0 31.0 26.0 71.0 53.0 52.0 43.9 26.9 46.3 41.3
FLAN-T5 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 24.0 31.0 48.0 20.0 84.0 78.0 64.0 39.8 35.0 51.5 45.3
  + CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 23.9 39.0 46.0 28.0 73.0 57.0 55.0 41.8 31.3 49.5 43.9
  + FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 23.0 31.0 47.0 21.0 82.0 78.0 65.0 41.1 34.2 51.3 45.1
  + FS/CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 19.0 34.0 45.0 20.0 89.0 72.0 66.0 42.9 31.1 50.7 44.7
Mistral 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 5.0 4.0 7.0 4.0 2.0 1.0 51.0 41.6 31.8 18.0 25.3
  + CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 11.0 8.0 14.0 12.0 5.0 4.0 61.0 36.4 23.5 21.2 24.2
  + FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 12.0 4.0 7.0 7.0 23.0 23.0 57.0 40.3 41.9 25.6 32.2
  + FS/CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.1 8.0 34.4 - 13.0 11.0 9.0 9.0 9.0 68.0 42.0 50.0 37.9 27.0 31.8 31.6
${}^{\dagger}$ 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 40.0 42.0 31.0 13.0 68.0 50.0 54.0 40.4 44.6 47.6 45.7
  + CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 39.0 39.0 34.0 37.0 73.0 46.0 60.0 46.8 39.8 50.9 47.6
${}^{\dagger}$ 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 34.0 41.0 38.0 34.0 68.0 49.0 58.0 40.9 43.7 50.0 46.9
  + CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 37.0 49.0 34.0 38.0 70.0 55.0 58.0 47.5 38.5 52.5 48.4
${}^{\dagger}$ 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 30.0 37.0 38.0 37.0 68.0 64.0 53.0 45.0 52.8 51.0 50.6
  + CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 30.0 48.0 37.0 24.0 60.0 58.0 57.0 52.3 44.3 49.6 48.7
Table 4: Main results of TimeBench. Aligned models are under the zero-shot setting by default. Methods with † are base models without alignment, under few-shot setting, thus incomparable with other methods. Some F1 scores are omitted, please refer to the appendix for details. Best results in each group are bold, top-3 results are underlined.
@@ 5 Experimental Results
5.1 Main Results
We present the experimental results [^2] Please refer to Appendix A.3 for full results. of the aligned models in Table 4, covering the most popular LLMs. In most tasks, GPT-4 achieved the best performance, ranking first in 16 out of 19 evaluation metrics, surpassing the second-place GPT-3.5 by 14.7%. Nevertheless, it still exhibits a significant gap from human performance, with a 19% disparity. Among all open-source alignment models, Baichuan2 has achieved the best overall performance, reaching 85% performance of ChatGPT.
Results on Symbolic Temporal Reasoning
GPT-4 demonstrates strong symbolic reasoning performance, far surpassing ChatGPT and other models. While open-source models show poorer performance, with only Mistral-base achieving an overall score exceeding 50. In TimeX-NLI tasks, models other than GPT-4 perform poorly, barely surpassing a random baseline (33.3 for s1/s3, 50.0 for s2). This indicates that LLMs still have significant shortcomings in understanding abstract TimeX.
Results on Commonsense Temporal Reasoning
In temporal commonsense reasoning, GPT-4 still maintains a remarkable proximity to human-level performance. Among the open-source aligned models, Baichuan2 attains the best performance, securing the top three positions in four out of seven metrics. When extending to all open-source models, Mistral-base significantly outperforms others, surpassing the best open-source model by 6.4. Besides, we observe that chain-of-thought prompting has a negative impact on almost all models in commonsense reasoning. Furthermore, aligned models appear to exhibit varying degrees of performance degradation compared to the base model, with LLaMA2 and Mistral showing the most significant differences.
Results on Event Temporal Reasoning
GPT-4 remains the top performer in event temporal reasoning, though it is no longer markedly ahead of other models, and there is a notable disparity from human performance. Among all open-source models, LLaMA2-base achieved the best results, trailing GPT-3.5 by just 1.5%. Overall, open-source models show competitive performance in event temporal reasoning, narrowing the gap with closed-source models.
In TRACIE, which involves implied events, most models only marginally surpass the random baseline (50.0), indicating persistent challenges in understanding implicit events with current models. Additionally, there is a large performance gap comparing context-based reasoning (TimeQA, MenatQA) with fact-based reasoning (TempReason), indicating a substantial potential for improvement in event-time facts grounding in unstructured text. Furthermore, within fact-based reasoning, models exhibit a noticeable decline in event-event reasoning compared to event-time reasoning, indicating that complex multi-hop relations among events increased the difficulty of reasoning.
Figure 3: Performance difference between base and aligned models. Baichuan2 and LLaMA2 are aligned with SFT and RLHF. Vicuna, Mistral and ChatGLM3 are aligned with only SFT.
@@ 6 Analysis and Discussion
6.1 Alignment Degradation
Alignment Causes Performance Degradation
We notice that some aligned models exhibited performances inconsistent with their expectations, such as LLaMA2 and Mistral. We speculate that this is the negative impact brought about by the alignment phase. To investigate this, we conduct experiments on their corresponding base models in the few-shot setting, as shown in Table 4 (grey).
We visualize the performance differences between the base and alignment models in the few-shot scenario, as shown in Figure 3. Baichuan2-13B shows a slight decrease in performance after alignment by 0.7%, while LLaMA2-13B, Vicuna-1.5-13B and Mistral-7B exhibit catastrophic performance drops, decreasing 10.0%, 7.1% and 18.4%, respectively.
What leads to the degradation?
Combining experimental results with model output cases, we attribute this phenomenon to two reasons:
  • The training during the alignment phase has to some extent disrupted the original language modeling capability, causing the model to deviate from strictly adhering to demonstrations during in-context learning.
  • Excessive alignment of models towards conversation and harmlessness has led to a decrease in its usability, resulting in generating instruction-irrelevant content or refusing to follow instructions.
Figure 4: Performance gap with and without CoT prompts. The results are averaged from GPT-4, GPT-3.5, Baichuan2, LLaMA2/Vicuna-1.5 and Mistral.
6.2 Chain-of-Thought in Temporal Reasoning
Previous research has found that chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2023a) enables models to think step by step, facilitating the resolution of complex reasoning tasks such as mathematical and logical reasoning. How about applying cot prompting to temporal reasoning, will it bring consistent improvement? To investigate this, we select several popular models and analyze their performance affected when using chain-of-thought prompts across various task categories.
Chain-of-Thought inconsistently boosts, or even impairs performance
As illustrated in Figure 4, introducing zero-shot chain-of-thought prompting results in an overall performance decline of 6%, with even a 16% degradation in commonsense reasoning. In the few-shot scenario, chain-of-thought prompts also fail to yield comprehensive improvements, showing varied effects depending on the task type, resulting in a 13% improvement in symbolic reasoning but an 18% decrease in commonsense reasoning. In event temporal reasoning, cot prompting in both scenarios only resulted in minor impacts, with -1% and 1.7%, respectively. In summary, zero-shot chain-of-thought consistently hinders performance, whereas few-shot chain-of-thought demonstrates performance variations dependent on the nature of the task.
Impact of Chain-of-Thought across tasks
In order to explore the impact of CoT on various tasks thoroughly, we delve into the performance changes of each model across specific tasks within each category, as illustrated in Figure  5. In the zero-shot setting, open-source models achieve a slight improvement in event temporal reasoning with chain-of-thought prompting, while in other cases, they face performance degradation. While in the few-shot setting, almost all models exhibit significant improvement in symbolic temporal reasoning, with a concurrent prevalent decline in commonsense temporal reasoning.
Differing from the previous uniform trends observed within the same task category, event temporal reasoning exhibits both increases and decreases simultaneously. We attribute this to the nature of the datasets; both MenatQA and TempReason involve implicit reasoning, encompassing temporal factors and event-event relationships, requiring multi-hop reasoning for resolution. This aligns well with the step-by-step problem-solving nature of cot prompting, resulting in performance improvement. On the other hand, TimeQA only necessitates implicit reasoning for a subset of questions, making cot prompts less impactful. This is reflected in the marginal performance change with and without chain-of-thought prompting.
6.3 Impact of Model Size
We investigate the relationship between the performance and parameter size of base models with decoder architecture ranging from 6B to 13B. With the scaling up of the parameter size for the same model, there is a notable performance enhancement. When the parameter size expands from 7B to 13B, LLaMA2 and Baichuan2 show performance improvements of 14.2% and 11.1%, respectively. Under the same 13B parameter size, LLaMA2 slightly outperforms Baichuan2 by 1.6%. With around 7B parameter size, there is no significant difference in performance among LLaMA2, Baichuan2, ChatGLM3, while Mistral exhibits extremely strong performance, even surpassing LLaMA2-13B by as much as 4.5% with only 53% parameters.
6.4 Dive into Each Categories
@@
${}^{\dagger}$ 40.5 $\downarrow$ 51.8 $\uparrow$ 43.7 $\uparrow$ 46.2 $\uparrow$ 29.8 $\downarrow$ 42.5
${}^{\dagger}$ 18.9 $\downarrow$ 37.0 $\uparrow$ 27.1 $\uparrow$ 7.6 $\downarrow$ 25.5 $\downarrow$ 26.5
${}^{\dagger}$ 27.0 $\downarrow$ 44.4 $\uparrow$ 58.3 $\uparrow$ 38.5 $\downarrow$ 38.3 $\downarrow$ 42.5
@@
Table 5: Results in each temporal commonsense aspect under few-shot setting. Models with †  are base models. Red $\downarrow$ and Green $\uparrow$ represent the performance is lower or higher than its average.
LLM underperforms in (multi-hop) symbolic reasoning
Except for GPT-4, the performance of other models in symbol temporal reasoning is unsatisfactory. Our experimental results reveal a notable decline in models’ performance on the duration-conversion task (s3) compared to other atomic tasks (s1, s2, arith). Delving into the reasons, task s3 necessitates a two-step reasoning process, initially unifying time units and subsequently engaging in numerical comparison. In contrast, other atomic tasks can be completed with a single inference. To sum up, LLMs show unsatisfactory performance in symbol temporal reasoning, with a more pronounced decline when confronted with multi-step reasoning.
LLM’s mastery of temporal commonsense knowledge differs across aspects.
@@
We analyze models’ performance across various common-sense aspects, as shown in Table 5. We use the model’s average performance in common-sense tasks as a baseline, if the model outperforms its own average in a specific aspect, it suggests greater proficiency and vice versa. The findings show that models generally demonstrate good knowledge of event duration and frequency, while their grasp of event order and typical events is comparatively weaker.
LLM demonstrates poor implicit temporal reasoning abilities
Comparing explicit and implicit context-based reasoning (TimeQA-explicit v.s. others), we note a marked decrease in the model’s performance in implicit reasoning. Additionally, on TRACIE with numerous implied events, most models exceed only a random baseline (50.0). Even GPT-4 achieves a mere 66.4% accuracy. To summarize, LLM performs inadequately when dealing with different types of implicit reasoning, constraining its reasoning abilities in complex scenarios.
LLM is a good factual reasoner rather than factual extractor
@@
When humans engage in context-based temporal reasoning, it typically involves two steps: initially extracting time-fact pairs from the context and subsequently conducting fact-based reasoning. TempReason provides extracted facts for conducting fact-based reasoning. By comparing the model’s performance in context-based reasoning and fact-based reasoning, we can identify the bottlenecks in LLM’s event temporal reasoning. Our findings reveal that LLM excels in TempReason, signifying its strong capability in fact-based reasoning. However, LLM’s performance in context-based reasoning was significantly weaker than in the former. We attribute this gap to the model’s deficiency in factual extraction capabilities. This gap highlights limitations in context-based temporal reasoning.

### v2

@@ 4.1 Models
[…]
@@ 4.2 Implementation Details
We access proprietary models through Azure API 0613 version. For open-source models, we deploy them locally through FastAPI. We set the temperature to 0.0 for greedy decoding in all experiments. To improve answer extraction accuracy, we prompt models with trigger Therefore, the answer is before model outputs to deduce final answers.
@@ 5 Experimental Results
5.1 Few-shot Results
[…]
In symbolic temporal reasoning tasks, GPT-4 demonstrates exceptional performance. However, other models exhibit a significant decline in comparison to GPT-4. In commonsense temporal reasoning tasks, GPT4 lags behind humans by only 8.0%, indicating its powerful internal knowledge reservoir. With the model scale shrinking, its knowledge reservoir also decreases gradually, leading to a decline in performance. Notably, there is a significant gap of 25.2% between LLMs and humans in event temporal reasoning, which suggests that LLMs encounter major challenges in modeling intricate event-time relationships.
5.2 Zero-shot Results
Experimental results of alignment models under zero-shot settings are shown in Table 5. In zero-shot settings, GPT-4 and GPT-3.5 rank first and second, respectively, and they significantly outperform all open-source models by a large margin. It is noteworthy that open-source models exhibit a larger performance decline compared to proprietary models when transitioning from few-shot to zero-shot scenarios. GPT, Baichuan2 and LLaMA2 suffer drops of 5.6%, 14.6% and 27.2%, respectively. We attribute this performance decline to the quality of alignment. Restricted by their limited instruction-following capability, open-source models struggle to fully unleash their performance solely through instructions. Therefore, few-shot prompting is a better approach for stimulating their temporal reasoning abilities.
[…]
5.3 Chain-of-Thought in Temporal Reasoning
Previous research has found that chain-of-thought prompting can enhance the model’s reasoning ability (Wei et al., 2022; Kojima et al., 2022). We aim to explore the following questions: Does CoT prompting bring consistent improvement in temporal reasoning? Due to the diversity of temporal reasoning, the above question has not yet been definitively answered. To investigate this, we select several popular LLMs and analyze their performance affected by chain-of-thought prompting.
Chain-of-thought reasoning is not consistently effective.
As illustrated in Figure 2, introducing zero-shot CoT prompting results in consistent declines, with an overall decrease of 7.4%. In the few-shot scenario, CoT prompting also fails to yield consistent improvements, varying depending on the task. There is a 10.8% improvement in symbolic reasoning, while a significant decline of 15.2% in commonsense reasoning. In event temporal reasoning, there is a slight improvement of 1.3%. Next, we will conduct a more detailed analysis of the impact of CoT on specific tasks.
Impact of CoT prompting across tasks.
In order to explore the impact of CoT on various tasks thoroughly, we delve into the performance changes of each model across specific tasks within each category, as illustrated in Figure 3. In the zero-shot setting, open-source models achieve a slight improvement in event temporal reasoning with chain-of-thought prompting, while in other cases, they face performance degradation. While in the few-shot setting, almost all models exhibit significant improvement in symbolic temporal reasoning, with a concurrent prevalent decline in commonsense temporal reasoning. We attribute this to the knowledge sensitivity inherent in commonsense reasoning, where step-by-step reasoning cannot compensate for the lack of knowledge. In event temporal reasoning, improvements mainly stem from datasets involving implicit multi-step reasoning (MenatQA and TempReason), indicating that CoT is more effective for multi-hop questions. In summary, zero-shot CoT consistently has a negative impact on temporal reasoning. While in few-shot scenario, CoT has a positive impact on symbolic and complex tasks, while negatively affecting knowledge-sensitive tasks.
@@ 6 Analysis and Discussion
6.1 Scaling Effect of Model Size
Figure 4: Scaling effect of model size and overall temporal reasoning performance. The x-axis (model size) is shown in the log scale. Results show a log-linearity between parameter size and performance.
We investigate how the scale of models affects temporal reasoning capabilities. The trend is illustrated in Figure 4. As the model scale increases, there is a notable improvement in performance. When the parameter size expands from 7B to 13B, LLaMA2 and Baichuan2 show improvements of 13.0% and 10.5%, respectively. Furthermore, when LLaMA scales up to 70B, the trend of performance improvement continues without stopping. The overall improvement follows a log-linear relationship with scale. There are no significant performance differences among LLaMA2, Baichuan2, and ChatGLM3 under similar parameter specifications, while Mistral demonstrates impressive prowess, outperforming all other 13B models with nearly half the number of parameters.
6.2 Challenges in Temporal Reasoning
@@
[…]
@@
Table 6: Results in each temporal commonsense aspect under few-shot setting. Models with †  are base models. Red $\downarrow$ and Green $\uparrow$ represent the performance is lower or higher than its average performance. Metric is EM.
LLMs underperform in (multi-hop) symbolic reasoning
[…]
Mastery of commonsense knowledge varies in LLMs
@@
We analyze models’ performance across various commonsense aspects, as shown in Table 6. We regard the model’s average performance in commonsense reasoning tasks as the baseline. If the model outperforms the baseline in a specific aspect, it suggests greater proficiency in this type of knowledge, and vice versa. The findings indicate that LLMs generally demonstrate good knowledge of event duration and frequency. However, their comprehension of event order and typical events is relatively weaker The uneven mastery of commonsense knowledge significantly affects the model’s reasoning performance, especially when dealing with complex questions that involve multiple types of knowledge. Retrieval-augmented reasoning presents a promising avenue for mitigating the model’s knowledge scarcity.
LLMs exhibit poor implicit temporal reasoning capabilities.
When comparing explicit and implicit event temporal reasoning, specifically TimeQA-explicit versus others, we observe a significant performance decrease in implicit reasoning. Additionally, on TRACIE with numerous implied events, most models only surpass a random baseline (50.0). Even GPT-4, despite its advanced capabilities, achieves only a 66.4% accuracy, suggesting that the LLM struggles with modeling implicit temporal relationships. We consider it helpful to explicitly model the temporal relationships between events and time expressions, for instance constructing timelines or temporal graphs.
LLMs are good factual reasoners rather than factual extractors
@@
When humans engage in temporal reasoning, it generally involves two steps: first, extracting time-fact pairs from the context, and then performing fact-based reasoning. TempReason provides extracted facts for conducting fact-based reasoning. By comparing the model’s performance in context-based (TimeQA) against fact-based (TempReason) reasoning, we identify the bottleneck in event temporal reasoning. LLMs excel in TempReason, which signifies their strong capability in fact-based reasoning. However, their performance in context-based reasoning is significantly weaker compared to their performance in fact-based reasoning. This implies that errors could arise during the extraction of time-sensitive facts from the context. We attribute this performance gap to the model’s deficiency in factual extraction capabilities Thus, we consider LLMs to be strong factual reasoners rather than factual extractors in event temporal reasoning.
Figure 5: Performance difference between base and alignment models under few-shot setting. Baichuan2 and LLaMA2 are aligned with SFT and RLHF. Vicuna, Mistral and ChatGLM3 are aligned with only SFT.
6.3 Alignment Impairs Temporal Reasoning
In the experiments mentioned earlier (Table 5), we observe a sharp decline in zero-shot performance of alignment models. To investigate whether alignment is the cause of the decline in temporal reasoning, we conducted experiments on alignment models under few-shot settings. Figure 5 illustrates the overall performance decline after alignment. With the exception of Baichuan2, all other models are severely impaired, experiencing a significant drop of up to 22%. Through manual analysis of error cases, we have summarized two reasons: (1) Alignment reduces the model’s usability, causing it to tend towards refusal to answer when confronted with knowledge-sensitive questions. (2) Alignment damages the model’s in-context learning capability, resulting in situations where the model deviates from the demonstrations. Furthermore, we believe that the lack of temporal reasoning-related training data in alignment exacerbates this issue, leading to disparities between different reasoning capabilities, such as mathematical and temporal reasoning.
Figure 6: Error analysis for Symbolic, Commonsense, and Event Temporal. We select 100 test samples from each subtask for GPT-4, GPT-3.5 and LLaMa2-base70b.
6.4 Error Analysis
We manually analyze 100 predictions by GPT-4, GPT-3.5 and LLaMa2-base70b from each subtask. The visualization of errors is shown in Figure  6.
Symbolic Reasoning
We categorize symbolic reasoning errors into five groups: (a) Expression: The model provides an incorrect time calculation expression. (b) Computation: The model provides the correct time calculation expression, but there is a calculation error. (c) Conversion: The model has an error in the conversion of time units. (d) Comparison: The model has an error when comparing two time-expressions (or intervals). (e) Combination: The model encounters errors in the combination of multiple above operations. LLMs exhibit numerous computation, conversion, and comparison errors, which suggests a substantial deficiency in their understanding of fundamental temporal expressions. Additionally, a higher frequency of errors is observed in combination questions, highlighting that multi-step reasoning continues to be a significant challenge for current models
Commonsense Reasoning
We categorize the errors of commonsense reasoning into two groups: (a) No Answer: The model fails to provide a final answer. (b) Reasoning Error: The model encounters reasoning errors, which can be subdivided into five types of knowledge-related errors. We observe that GPT series models have a higher No Answer rate, while LLaMA is always able to provide answers. This discrepancy can be attributed to two factors: firstly, the models may lack the necessary commonsense knowledge to formulate an answer; secondly, the preference alignment mechanism may prompt the model to abstain from answering when confronted with questions outside its knowledge scope. Integration of retrieval can alleviate the problem of knowledge scarcity to a certain degree.
Event Temporal Reasoning
We categorize the errors of event temporal reasoning into four groups: (a) No Answer: The model is unable to find the answer in the context. (b) Reasoning Error: The model encounters reasoning errors. (c) Hallucination: The model’s prediction does not exist in the context, known as hallucination reasoning. (d) Metric: The model’s prediction is correct, but the metric is limited by the evaluation criteria. It can be observed that, except for reasoning errors, failures to provide answers account for approximately 30%, indicating that models still have flaws in grounding temporal facts from context. Additionally, models occasionally experience hallucination phenomena, leading to erroneous reasoning.